A firm’s stock is selling for $134. They just paid a $4 dividend and dividends are expected to grow at 8% per year. 

.1. What is the required return?

.2. What is the dividend yield?

.3. What is the capital gains yield?



1
Expert's answer
2021-04-14T14:58:30-0400

To get the required return

We use the dividend growth model

g=growth rate per year

re=return

Div=Dividends paid

= [Div(1+g)]/[ (re-g)]

134=4(1+ 0.08)/[(re-0.08)

134re - 10.72 = 4.32

134re = 10.72 + 4.32

re = 11.2238806

= 11.2239%

.2. What is the dividend yield?

To get the dividend yield: It is the return from dividends


=4(1.08) / 134 = 3.2238805%


.3. What is the capital gains yield?

this is the return from change in stock price

In this case =g=8%